Can you please assist me with the following equation: y=2x+6
y=8x-34
I've been trying for more than an hour to solve this and I'm still unable to do so. Any assistance that you can provide will be greatly appreciated. Also, can you please show the steps? Thank you.

Answer by ewatrrr(24785)

y=2x+6 and y=8x-34
Therefore...
2x+ 6 = 8x - 34
40= 6x
40/6 = x
x = 20/3